A student wants to establish a standing wave on a wire 1.8 m long clamped at both ends. If the wave speed is 72 m/s, what is the minimum frequency she should apply to set up standing waves?


20 Hz


12 Hz


24 Hz


15 Hz


18 Hz


1
Expert's answer
2021-05-23T18:55:26-0400

"\frac{2l}{v}=nT\implies \frac{2l}{\lambda}=n,"

"\nu=\frac 1T=\frac{v}{\lambda}=\frac{v}{2l}n,"

"\nu_{min}=\nu_{n=1}=\frac{v}{2l}=\frac{72}{2\cdot 1.8}=20~Hz."